OpenStudy (anonymous):

volume is 7599 m^3 and height is 20m what is radius?

OpenStudy (anonymous):

This is a right-circular-cylinder?

OpenStudy (anonymous):

V = area of circle * height = 7599 V = pi*r^2 * 20 = 7599 7599 / 20 = pi*r^2 \[\sqrt{\frac{7599}{20*pi}} = r\]

OpenStudy (anonymous):

we ignore the negative answer because radius will be positive

OpenStudy (anonymous):

so far so good

OpenStudy (anonymous):

\[r \approx 10.997\], but if this is homework, your book may say \[r \approx 11\]by rounding up

OpenStudy (anonymous):

gotcha, thank you

OpenStudy (anonymous):

yw